OpenStudy (technodynamic):

Let h(x)=1/(x^2) a) Find h(x+8)= 1/(x+8)^2 b) Find h(x)+8= c) Find h(x)+h(8)=

OpenStudy (anonymous):

h(x) + 8 = 1/(x^2) + 8 = (1 + 8(x^2))/(x^2) h(x) + h(8) = 1/(x^2) + 1/64 = (64 + x^2)/(64x^2)